OpenStudy (anonymous):

I will medal best answer!!! Find the unknown number. Five added to 2 times a number is equal to 7 less than 4 times the number. 6 3 –6 –3

OpenStudy (anonymous):

5+2*x=-7+4*x then rearrange: 12 = 2*x solve: x = 6

OpenStudy (anonymous):

Let the number be x. Then, convert the words into math: \[ 5 + 2x = 4x - 7 \implies 2x = 12 \implies x = \boxed {6}. \]

OpenStudy (anonymous):

6